Find coordinates x]B of relative to the basis B = {v1, v2}, where
Answer #1

e[4= * » [12] [% = ] (-not an2 =-1) x3 -32, t602=-3 ini -442 = g 31, 412 = 9 - 202 = 6 - N2 = 3 ni = 2n2t1 = 2(3)+1 3 7. ol t